OpenStudy (anonymous):

Math Humor:

OpenStudy (anonymous):

In the words of Yoda. \[ \pi, \hspace{2mm}\sqrt{-1}\hspace{4mm} \sqrt{64} \] Delicious it was. Get it ? HA HA

jimthompson5910 (jim_thompson5910):

funny that is...

OpenStudy (across):

I have a shirt that reads:\[\text{Math is}\int e^{xy}\]

OpenStudy (anonymous):

Wish you could give more than one point :-(

OpenStudy (anonymous):

I don't get it....:(

OpenStudy (anonymous):

Math Humor. In the words of Yoda. \( \pi \) = pie \(\sqrt{-1} \) = i \( \sqrt{64} = 8\) Delicious it was.

OpenStudy (anonymous):

Ahhh....how dumb of me...

OpenStudy (anonymous):

hey, @across, what's yours?

OpenStudy (anonymous):

across's says Math is sexy

jimthompson5910 (jim_thompson5910):

or how about \[\large \sqrt{-1} \ \ 8 \ \text{ your } \pi ... \frac{\text{get}}{\text{it}}\]

OpenStudy (anonymous):

lol NO..... not my last piece lol

jimthompson5910 (jim_thompson5910):

lol \(\frac{\sqrt{64}}{4}\) bad

OpenStudy (anonymous):

yum

OpenStudy (anonymous):

\( \sqrt{-1} \hspace{4mm} \sqrt{64} \hspace{4mm} \sum \hspace{4mm} \pi \) i 8 sum pi

OpenStudy (anonymous):

Oh, how I wish I could give more than 1 medal per question...

jimthompson5910 (jim_thompson5910):

\( \sqrt{-1} \hspace{4mm} \sqrt{64} \hspace{4mm} \sum \hspace{4mm} \pi = \infty\)

OpenStudy (anonymous):

You like PIE lol

OpenStudy (anonymous):

I say you bump this to the top so everybody can see it.